OpenStudy (anonymous):

root:6:(a^(4)) Write in radical notation.

OpenStudy (anonymous):

\[\sqrt[6]{a^{4}}\] is that it ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

write in radical notation

OpenStudy (anonymous):

can somoene help

OpenStudy (anonymous):

I think that is radical notation...

OpenStudy (anonymous):

\[(a ^{2}b ^{2}) 1/5\]

OpenStudy (anonymous):

You can simplify it somewhat if you like..

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[\sqrt[6]{a^4} = \sqrt[3]{a^2}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Do you understand why that works?

OpenStudy (anonymous):

kind of.

OpenStudy (anonymous):

I think

OpenStudy (anonymous):

\[or :\sqrt{a ^{3}}\]

OpenStudy (anonymous):

no I was wrong